OpenStudy (anonymous):

@amistre64 Solve the initial value problem below: (t^2 + 2t)dx/dt = 2x + 5 with t, x>0 and x(1) = 1

OpenStudy (anonymous):

Why do you keep reposting it it was solved for x \[x=\frac{7\sqrt3}{2}(\frac{t}{t+2})-5\]

OpenStudy (anonymous):

that isn't correct.. :/

OpenStudy (anonymous):

Sorry correction \[x=\frac{7}{3√2}\sqrt{(\frac{t}{t+2})}−5\]

OpenStudy (anonymous):

ahhh, let me try that.

OpenStudy (anonymous):

that's not right either..

OpenStudy (amistre64):

system froze ...

OpenStudy (anonymous):

uh oh.. not good.

OpenStudy (amistre64):

make sure your question is posted correctly; no typos please

OpenStudy (anonymous):

ok I'll screen shot it.

OpenStudy (anonymous):

IT is \[\frac{7\sqrt3}{2}\] it was an typing error

OpenStudy (anonymous):

it still says that is incorrect

OpenStudy (anonymous):

aw it also -5/2 I am sorry i usually mess up the numbers

OpenStudy (anonymous):

No worriest at all. let me try that.

OpenStudy (anonymous):

by 5/2 do you mean.. ( 7/(3sqrt(2)) * sqrt( t / (t+2)) -5/2)?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

didn't work either... :/

OpenStudy (amistre64):

\[(2x + 5)^{-1}dx = (t^2 + 2t)^{-1} dt \] \[\frac 12 ln(2x + 5) = \int (t^2 + 2t)^{-1} dt \] gotta partial that up i think 1 A B ----- = -- + --- t(t+2) t t+2 1 = A(t+2) + Bt ; when t=0; A=1/2 when t=-2 B=-1/2 \[\frac 12 ln(2x + 5) = \int \frac{1/2}{t}-\frac{1/2}{t+2} dt \] \[\frac 12 ln(2x + 5) = \frac{1}{2}ln(t)-\frac{1}{2}ln(t+2) +C \] \[ln(2x + 5) = ln(t)-ln(t+2) +C \] \[ln(2x + 5) = ln(\frac{t}{t+2}) +C \] \[2x + 5 = \frac{Ct}{t+2} \] maybe

OpenStudy (anonymous):

Aw i messed up the lhs @amistre64 is right

OpenStudy (amistre64):

so if its good; x(1) = 1 \[7 = \frac{C}{3}\] C = 21

OpenStudy (amistre64):

\[x = \frac{21t}{2(t+2)}-\frac52\] but id have to check the wolf

OpenStudy (anonymous):

That sir, is correct. you rock. NotSObright, thanks for your help alot too bud.

OpenStudy (amistre64):

yay!!